Diffusion Morning report 11/15/10

Ted Berg tberg at stanford.edu
Mon Nov 15 06:32:48 PST 2010


Good Morning All,
     Looking OK this morning.Thermco Polys have leak issues. Epi being 
baked after Dome change-needs test. Ebeam still has computer issues. LTO 
going down tomorrow for pull and clean. No other issues at this time.

					
	Date: 	11/15/10 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*yellow* 	*intermittent mfs alarms and n2 flow issues* 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*20.1/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*12.8/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*16.86/30* 	** 	**
Tylan 11-LTO 	*up* 	*going down tomorrow for pull & clean* 
*31.01/30* 	** 	**
Tylan 12-FGA 	*up* 	*
* 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*yellow* 	*leak issues* 	** 	** 	**
Thermco Poly2 	*up* 	*leak issues* 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*yellow* 	*testing* 	** 	** 	**
AMAT Epi 	*down* 	*dome replaced- baking* 	** 	** 	**
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*computer issues* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 633 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20101115/55fe26bf/attachment.html>


More information about the diffusion mailing list